# Kac–Moody Symmetric Spaces: arbitary symmetrizable complex or almost split real type

### v3

@@ 6.2 A Local to Global Approach
Here $\kappa$ and $\iota$ denotes canonical inclusion. Further, any local automorphism extends to a global automorphism if and only if $\operatorname{Aut}(\Gamma)=\operatorname{Aut}(\mathcal{W},S)$.
@@
The Chevalley involution $\Theta$ acts on the standard flat by inversion and so it corresponds to $\mathbb{Z}/2\mathbb{Z}$. Note that the semidirect product with $\Theta$ vanishes in the decomposition of the stabilizer in $\operatorname{Aut}_{\text{eff}}(G^{\dagger})$ as one considers the action on $A^{\dagger}K^{\dagger}$ and by the proof of [FHHK20, Proposition 6.5] it is known that the Chevalley involution commutes with $\operatorname{Aut}(\Gamma)$ and it can be concluded that $\Theta$ also commutes with the inner automorphisms. In fact, since one considers the stabilizers of the standard flat $A^{\dagger}K^{\dagger}$, one know that $\Theta$ commutes with the diagram automorphism and normalizes the left multiplication. If we denote the left multiplication with $g\in\operatorname{Stab}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})$ by $l(g)$, then the following is true
@@
After this, it is left to analyze the first factor in the decomposition of $\operatorname{Aut}(X^{\dagger})$. Note that the inner automorphism of $\operatorname{Aut}_{\text{eff}}(G^{\dagger})$ corresponds to $G^{\dagger}/C_{K^{\dagger}}(G^{\dagger})$, cp. 5.5. Further, by 6.5 one can consider the stabilizer and fixator of $A^{\dagger}K^{\dagger}$ as subgroups of $G^{\dagger}$:
@@ 7 Comparing Symmetric Spaces
In the previous chapters we established a symmetric space associated to Kac–Moody groups over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and with respect to arbitrary generalized Cartan matrices. Now we can give an summary and compare it with the symmetric space developed in [FHHK20], which was set up for non-affine Kac–Moody groups over $\mathbb{R}$.
@@
This larger torus turns out to be a problem, because for the analysis of the global and local action of the Kac–Moody group on the corresponding symmetric space it is important that the two automorphism groups are related (see 5.5). The reason for this problem is Caprace’s theorem: It is the basis for understanding the automorphisms of a particular type of Kac–Moody group. But it only describes the automorphisms of a Kac–Moody group generated by its root groups, see [Cap09, Theorem 4.2]. On the extended torus one essentially asks for the automorphism group of $(\mathbb{R},+)$. This group is too big to classify, so it is not possible to say anything about the automorphisms on the torus outside of $G^{\dagger}$.
@@ 8.1 Real Forms of Kac–Moody Groups
  • A $\mathbb{R}$-functor $F$ is called a $\mathbb{R}$-form of a $\mathbb{C}$-functor $H$, if there is a functorial isomorphism $F\mid_{\mathbb{C}}\cong H$.
Clearly is the Tits functor $\mathcal{G}$ an example of an $\mathbb{R}$-form of itself.
Let $F$ be a $\mathbb{R}$-form of a $\mathbb{C}$-functor $H$. Now, we want to define an action of the Galois group $\Gamma\coloneqq\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ on $H(\mathbb{C})$ using the real form $F$. Therefore denote by $R_{\mathbb{R}}$ a $\mathbb{R}$-algebra and by $R_{\mathbb{C}}=R_{\mathbb{R}}\otimes_{\mathbb{R}}\mathbb{C}$ the corresponding $\mathbb{C}$-algebra given by scalar extension. Then for all $\sigma\in\Gamma$ the map
@@
is an isomorphism of $\mathbb{R}$-algebras. By the fact, that the the real form $F$ equals the $\mathbb{C}$-functor $H$ over $\mathbb{C}$, the isomorphism $1_{R_{\mathbb{R}}}\otimes\sigma$ induces an automorphism $f_{\sigma}$ on $H(R_{\mathbb{C}})$. This leads to the following commutative diagram.
@@
  • For each subfield $\mathbb{K}$ of $\mathbb{C}$ is $\mathcal{F}(\mathbb{K})$ a subgroup of $\mathcal{G}(\mathbb{C})=G$.
@@
From [Rém02, Proposition 11.2.2] it follows that the Galois action on a Kac–Moody group $G$ stabilizes the set of tori in $G$ and that for every Galois action $\sigma$ there is an element $g\in G$, so that one can rectify the action in the sense that the standard torus is fixed. In detail, this means that the action
@@
    • the action $\overline{\sigma}$ respects the decomposition into $Q$-homogeneous components of $\mathcal{U}_{\mathbb{R}}$ and the induced permutation on $Q$ satisfies the following homogeneity condition:
      For all $\alpha\in\Delta^{\mathrm{re}}$ and for all $n\in\mathbb{N}$ holds: $\overline{\sigma}(n\alpha)=n\overline{\sigma}(\alpha)$.
@@
Note that by [Rém02, Lemma 11.3.2] one can rectify a Galois automorphism $\sigma$ in such a way that the sign of the Borel subgroup is preserved. Moreover, this rectification also stabilizes the standard torus $T$ and is well-defined up to $T$. Thus, this rectification is just a special case of the rectified Galois automorphism we defined above. In particular, this action induces an action on the Weyl group $\mathcal{W}$ and stabilizes the set of generators $S$, see [Rém02, Proposition 11.3.2]. Since we are only interested in almost split Kac–Moody $\mathbb{R}$ groups and for simplicity, we denote this rectification by $\overline{\sigma}$. For more details about almost split Kac–Moody $\mathbb{R}$-groups see [Rém02, 11.3.1].
Before we discuss Galois descent, we give useful properties of the rectified Galois action. Let $G$ be a Kac–Moody group over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is indecomposable and symmetrizable. Further let $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$ be a algebraic $\mathbb{R}$-form of $G$. Then any rectified Galois automorphism $\overline{\sigma}$ has the following properties:
@@
For details we refer to [Rém02, Proposition 11.2.5].
@@ 8.2 The Galois Descent
In linear algebra it is well known that a real form of a complex vector space can be identified with the fixed point set of an action given via Galois automorphism. The same idea uses Rémy to develop further the theory of almost split Kac–Moody $\mathbb{R}$-groups.
In [Rém02, Chapter 12], he uses the fact that one can associated to a Kac–Moody group a twin building on which he can also define a Galois action which is compatible with the Galois automorphism action on the group. In detail, he uses a geometric realization to study the Galois action and for example to check if there exists at all fixed points, for the realization see [Rém02, Chapter 5] and for the action on the realization of the twin building see [Rém02, 12.1.2].
@@
To make the following terms well-defined, we need to define a triple $(A,C,-C)$, where $A$ is an apartment and $C,-C$ are opposite chambers. This is similar to defining a standard torus and defines the geometric realization. In our context, it is only necessary that we obtain the following when applying the theory of [Rém02, Chapter 12]:
  • The group of rational points possess a root system, which is not necessarily reduced. A real $\mathbb{R}$-root is a restricted root in the context of the geometric realization. To be precise, a root (there it is called half-apartment) in the geometric realization is given by a hyperplane on which a reflection is defined. In the process of doing the Galois descent, one restricts the hyperplane to the fixed point set. For more details see [Rém02, Chapter 5]. There the geometric realization and the relative terminology is defined.
@@ 8.3 Relative Kac–Peterson Topology
Even the torus $Z(\mathbb{R})\leq G(\mathbb{R})$, defined in [Rém02, Chapter 12.5], is a subgroup of the torus of the surrounding split Kac–Moody group and splits over $\mathbb{R}$, see [Rém02, 12.5.1] and [Rém02, 12.5.3]. This means that, that $Z(\mathbb{R})$ also carries a Lie group topology.
@@
In the proof of [HK23, Porposition B.3] bounded subgroups are used to define a topology on a split Kac–Moody group $G$ which euqals the Kac–Peterson topology. This can also be done for rational points of an almost split Kac–Moody $\mathbb{R}$-group $G(\mathbb{R})$. To define the notion of a restricted subgroup in $G(\mathbb{R})$, we recall what a parabolic subgroup is in the context of a split Kac–Moody group $G$ over $\mathbb{C}$, see [Mar18, p.150]:
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix. Equip the standard torus with the Lie topology induced by the exponential map, see 3.7. Then the Kac–Peterson topology $\tau_{\text{KP}}$ equals the maximal bounded subgroup topology $\tau_{\text{MB}}$.
@@
We follow the strategy of the proof given for [HK23, Proposition B.3]. Since the cited proposition is given for centered Kac–Moody groups and in general a Kac–Moody group $G$ can be decomposed in $G=TG^{\dagger}$, see 3.10, where $G^{\dagger}$ is a centered Kac–Moody group, see 3.13, we has to take a closer look at the tours.
@@
and carries a Lie group topology, see 3.7. Thus the torus embeds as a closed subgroup into a maximal bounded subgroup and by the closed-subgroup theorem ([HN12, Theorem 9.3.7]) follows that the subspace topology of the embedded torus equals its Lie group topology.
@@
From the closed-subgroup Theorem (cf. [HN12, Theorem 9.3.7]) and the fact that a relative rank one subgroup is a closed subgroup in a maximal bounded subgroup of $\operatorname{MBS}(\mathbb{C})$ because the Galois action on $G$ is continuous, it follows that a relative rank one subgroup is an embedded Lie group. Moreover, the Lie group topology coincides with the subspace topology from $\operatorname{MBS}(\mathbb{C})$. Note that the subspace topology induced by $\tau_{\text{KP}}$ on a maximal bounded group of $\operatorname{MBS}(\mathbb{C})$ is equal to the Lie group topology, see the proof of 8.8.
@@
The universal topology $\tau_{\text{Prod}}$, defined in [HKM13, Definition 7.19] using colomits, is equal to the Kac–Peterson topology (cp. [HKM13, Proposition 7.21]) and by transport of structure, it is a $k_{\omega}$ topology, for details see Section 3.6 or [FT77]. The authors of the article [GGH10, Corollary 6.7] states that $G(\mathbb{R})$ is $k_{\omega}$ with respect to the induced subspace topology coming from the surrounding split Kac–Moody group.
@@
    Denote by $(\operatorname{MBS}(\mathbb{R})_{i})_{i\in I}$ the family of maximal bounded $\mathbb{R}$-subgroups in $G(\mathbb{R})$ and consider the topological spaces $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$ and $(G(\mathbb{R}),\tau_{\text{MB}}^{\operatorname{as}})$. A relative rank one subgroup embeds as a Lie subgroup into one of the $\operatorname{MBS}(\mathbb{R})_{i}$. Moreover the continuous embeddings of the relative rank one subgroups into $G(\mathbb{R})$ provide fewer conditions than the embeddings of the maximal bounded subgroups. It follows that the map from $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$ to $(G(\mathbb{R}),\tau_{\text{MB}}^{\operatorname{as}})$ is continuous, i.e. $\tau_{\text{MB}}^{\operatorname{as}}\subseteq\tau_{\text{KP}}^{\operatorname{as}}.$
    For the other direction consider the maximal bounded subgroups in the topological group $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$. Every $\operatorname{MBS}(\mathbb{R})_{i}$ carries the Lie group topology and the embedded relative rank one subgroups carry the induced Lie subspace topology. Thus one deduces that the embedding of $\operatorname{MBS}(\mathbb{R})_{i}$ into $G(\mathbb{R})$ with respect to the topology $\tau_{\text{KP}}^{\operatorname{as}}$ is continuous. To be more precise, the topology $\tau_{\text{KP}}^{\operatorname{as}}$ comes from the embedding of relative rank one subgroups equipped with the Lie topology and by the closed-subgroup theorem (cf. [HN12, Theorem 9.3.7]) the maximal bounded subgroups carries the same Lie group topology. This means that the topology $\tau_{\text{KP}}^{\operatorname{as}}$ is coarser than $\tau_{\text{MB}}^{\operatorname{as}}$.

### v4

@@ 6.2 A Local to Global Approach
Here $\kappa$ and $\iota$ denote canonical inclusion. Further, any local automorphism extends to a global automorphism if and only if $\operatorname{Aut}(\Gamma)=\operatorname{Aut}(\mathcal{W},S)$.
@@
The Chevalley involution $\Theta$ acts on the standard flat by inversion and so it corresponds to $\mathbb{Z}/2\mathbb{Z}$. Note that the semidirect product with $\Theta$ vanishes in the decomposition of the stabilizer in $\operatorname{Aut}_{\text{eff}}(G^{\dagger})$ as one considers the action on $A^{\dagger}K^{\dagger}$ and by the proof of [FHHK20, Proposition 6.5] it is known that the Chevalley involution commutes with $\operatorname{Aut}(\Gamma)$ and it can be concluded that $\Theta$ also commutes with the inner automorphisms. In fact, since one considers the stabilizers of the standard flat $A^{\dagger}K^{\dagger}$, one knows that $\Theta$ commutes with the diagram automorphism and normalizes the left multiplication. If we denote the left multiplication with $g\in\operatorname{Stab}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})$ by $l(g)$, then the following is true
@@
After this, it remains to analyze the first factor in the decomposition of $\operatorname{Aut}(X^{\dagger})$. Note that the inner automorphism of $\operatorname{Aut}_{\text{eff}}(G^{\dagger})$ corresponds to $G^{\dagger}/C_{K^{\dagger}}(G^{\dagger})$, cf. 5.5. Further, by 6.5 one can consider the stabilizer and fixator of $A^{\dagger}K^{\dagger}$ as subgroups of $G^{\dagger}$:
@@ 7 Comparing Symmetric Spaces
In the preceeding sections we constructed symmetric spaces associated to Kac–Moody groups over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and with respect to arbitrary symmetrizable generalized Cartan matrices. Now we can give an summary and compare it with the symmetric space developed in [FHHK20], which was set up for non-affine Kac–Moody groups over $\mathbb{R}$.
@@
This larger torus turns out to be a problem, because for the analysis of the global and local action of the Kac–Moody group on the corresponding symmetric space it is important that the two automorphism groups are related (see 5.5). The reason for this problem is Caprace’s theorem: It is the basis for understanding the automorphisms of a particular type of Kac–Moody group. But it only describes the automorphisms of a Kac–Moody group generated by its root groups, see [Cap09, Theorem 4.2]. On the extended torus one essentially asks for the automorphism group of $(\mathbb{R},+)$. This group is too big to be classified, so it is not possible to say anything about the automorphisms on the torus outside of $G^{\dagger}$.
@@ 8.1 Real Forms of Kac–Moody Groups
  • An $\mathbb{R}$-functor $F$ is called a $\mathbb{R}$-form of a $\mathbb{C}$-functor $H$, if there is a functorial isomorphism $F\mid_{\mathbb{C}}\cong H$, i.e. the functor $F$ applied to the field $\mathbb{C}$ is functorial isomorphic to the functor $H$.
The Tits functor $\mathcal{G}$ is clearly an example of an $\mathbb{R}$-form of itself.
Let $F$ be an $\mathbb{R}$-form of a $\mathbb{C}$-functor $H$. Now, we want to define an action of the Galois group $\Gamma\coloneqq\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ on $H(\mathbb{C})$ using the real form $F$. Therefore denote by $R_{\mathbb{R}}$ an $\mathbb{R}$-algebra and by $R_{\mathbb{C}}=R_{\mathbb{R}}\otimes_{\mathbb{R}}\mathbb{C}$ the corresponding $\mathbb{C}$-algebra given by scalar extension. Then for all $\sigma\in\Gamma$ the map
@@
is an isomorphism of $\mathbb{R}$-algebras. By the fact, that the real form $F$ equals the $\mathbb{C}$-functor $H$ over $\mathbb{C}$, the isomorphism $1_{R_{\mathbb{R}}}\otimes\sigma$ induces an automorphism $f_{\sigma}$ on $H(R_{\mathbb{C}})$. This leads to the following commutative diagram.
@@
  • For each subfield $\mathbb{K}$ of $\mathbb{C}$, $\mathcal{F}(\mathbb{K})$ is a subgroup of $\mathcal{G}(\mathbb{C})=G$.
@@
From [Rém02, Proposition 11.2.2] follows that the Galois action on a Kac–Moody group $G$ stabilizes the set of tori in $G$ and that for every Galois action $\sigma$ there is an element $g\in G$, so that one can rectify the action in the sense that the standard torus is fixed. Denote with $c_{g}(x)$ the map that conjugates an element $x\in G$ with $g$. Then we can state that the action
@@
Recall that $\mathcal{U}_{\mathbb{R}}$ is the universal enveloping algebra and therefore satisfies the same grading as the roots of the underlying Kac–Moody algebra with respect to $Q$. For details of the grading of the universal enveloping algebra see [Mar18, Chapter 3.1] and for the grading of the roots with respect to $Q$ see Equation 2.
@@
    • the action $\overline{\sigma}$ respects the decomposition into $Q$-homogeneous components of $\mathcal{U}_{\mathbb{R}}$ (i.e. it respects the $Q$-gradation) and the induced permutation on $Q$ satisfies the following homogeneity condition:
      For all $\alpha\in\Delta^{\mathrm{re}}$ and for all $n\in\mathbb{N}$ we have $\overline{\sigma}(n\alpha)=n\overline{\sigma}(\alpha)$.
@@
Note that by [Rém02, Lemma 11.3.2] one can rectify a Galois automorphism $\sigma$ in such a way that the sign of the Borel subgroup is preserved. Moreover, this rectification also stabilizes the standard torus $T$ and is well-defined up to $T$. Thus, this rectification is just a special case of the rectified Galois automorphism we defined above. In particular, this action induces an action on the Weyl group $\mathcal{W}$ and stabilizes the set of generators $S$, see [Rém02, Proposition 11.3.2]. Since we are only interested in almost split Kac–Moody $\mathbb{R}$-groups and for simplicity, we denote this rectification by $\overline{\sigma}$. For more details about almost split Kac–Moody $\mathbb{R}$-groups see [Rém02, 11.3.1].
Before we discuss Galois descent, we give useful properties of the rectified Galois action. Let $G$ be a Kac–Moody group over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is indecomposable and symmetrizable. In addition, let $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$ be an algebraic $\mathbb{R}$-form of $G$. Then any rectified Galois automorphism $\overline{\sigma}$ has the following properties:
@@
For details, we refer to [Rém02, Proposition 11.2.5].
@@ 8.2 The Galois Descent
In linear algebra it is well known that a real form of a complex vector space can be identified with the fixed point set of an action given via Galois automorphism. Rémy uses the same idea to develop further the theory of almost split Kac–Moody $\mathbb{R}$-groups.
In [Rém02, Chapter 12], he uses the fact that one can associate to a Kac–Moody group a twin building on which he can also define a Galois action which is compatible with the Galois automorphism action on the group. In detail, he uses a geometric realization to study the Galois action and for example to check if there exists at all fixed points, for the realization see [Rém02, Chapter 5] and for the action on the realization of the twin building see [Rém02, 12.1.2].
@@
To make the following terms well-defined, we need to define a triple $(A,C,-C)$, where $A$ is an apartment and $C,-C$ are opposite chambers in $A$. This is similar to defining a standard torus and defines the geometric realization. In our context, it is only necessary that we obtain the following when applying the theory of [Rém02, Chapter 12]:
  • The group of rational points possesses a root system, which is not necessarily reduced. A real $\mathbb{R}$-root is a restricted root in the context of the geometric realization. To be precise, a root (there it is called half-apartment) in the geometric realization is given by a hyperplane on which a reflection is defined. In the process of doing the Galois descent, one restricts the hyperplane to the fixed point set. For more details see [Rém02, Chapter 5]. There the geometric realization and the relative terminology is defined.
@@ 8.3 Relative Kac–Peterson Topology
Even the torus $Z(\mathbb{R})\leq G(\mathbb{R})$, defined in [Rém02, Chapter 12.5], is a subgroup of the torus of the surrounding split Kac–Moody group and splits over $\mathbb{R}$, see [Rém02, 12.5.1] and [Rém02, 12.5.3]. This means, that $Z(\mathbb{R})$ also carries a Lie group topology.
@@
In the proof of [HK23, Proposition B.3] bounded subgroups are used to define a topology on a split Kac–Moody group $G$ which equals the Kac–Peterson topology. This can also be done for rational points of an almost split Kac–Moody $\mathbb{R}$-group $G(\mathbb{R})$. To define the notion of a restricted subgroup in $G(\mathbb{R})$, we recall what is a parabolic subgroup in the context of a split Kac–Moody group $G$ over $\mathbb{C}$, see [Mar18, p.150]:
@@
Let $G$ be a split Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix. Equip the standard torus with the Lie topology (see discussion after 3.6 or 3.7), then the Kac–Peterson topology $\tau_{\text{KP}}$ equals the maximal bounded subgroup topology $\tau_{\text{MB}}$.
@@
We follow the strategy of the proof given for [HK23, Proposition B.3]. Since the cited proposition is given for centered Kac–Moody groups and in general a Kac–Moody group $G$ can be decomposed in $G=TG^{\dagger}$, see 3.10, where $G^{\dagger}$ is a centered Kac–Moody group, see 3.13, we have to take a closer look at the torus.
@@
and carries a Lie group topology, see discussion after 3.6 or 3.7. Thus the torus embeds as a closed subgroup into a maximal bounded subgroup and by the closed-subgroup theorem ([HN12, Theorem 9.3.7]) follows that the subspace topology of the embedded torus equals its Lie group topology.
@@
Recall that a topological group is called almost connected if the quotient of the group by the connected component of the neutral element is compact, see [Mar15].
@@
From the closed-subgroup Theorem (cf. [HN12, Theorem 9.3.7]) and the fact that a relative rank one subgroup is a closed subgroup in a maximal bounded subgroup of $\operatorname{MBS}(\mathbb{C})$ because the Galois action on $G$ is continuous, we conclude that a relative rank one subgroup is an embedded Lie group. Moreover, the Lie group topology coincides with the subspace topology from $\operatorname{MBS}(\mathbb{C})$. Note that the subspace topology induced by $\tau_{\text{KP}}$ on a maximal bounded group of $\operatorname{MBS}(\mathbb{C})$ is equal to the Lie group topology, see the proof of 8.8.
@@
The universal topology $\tau_{\text{Prod}}$, defined in [HKM13, Definition 7.19] using colimits, is equal to the Kac–Peterson topology (cf. [HKM13, Proposition 7.21]) and by transport of structure, it is a $k_{\omega}$ topology, for details see Section 3.6 or [FT77]. The authors of the article [GGH10, Corollary 6.7] state that $G(\mathbb{R})$ is $k_{\omega}$ with respect to the induced subspace topology coming from the surrounding split Kac–Moody group.
@@
    Denote by $(\operatorname{MBS}(\mathbb{R})_{i})_{i\in I}$ the family of maximal bounded $\mathbb{R}$-subgroups in $G(\mathbb{R})$ and consider the topological spaces $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$ and $(G(\mathbb{R}),\tau_{\text{MB}}^{\operatorname{as}})$. A relative rank one subgroup embeds as a Lie subgroup into one of the $\operatorname{MBS}(\mathbb{R})_{i}$. Moreover the continuous embeddings of the relative rank one subgroups into $G(\mathbb{R})$ provide fewer conditions than the embeddings of the maximal bounded subgroups. Hence the map from $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$ to $(G(\mathbb{R}),\tau_{\text{MB}}^{\operatorname{as}})$ is continuous, i.e. $\tau_{\text{MB}}^{\operatorname{as}}\subseteq\tau_{\text{KP}}^{\operatorname{as}}.$
    For the other direction consider the maximal bounded subgroups in the topological group $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$. Every $\operatorname{MBS}(\mathbb{R})_{i}$ carry the Lie group topology and the embedded relative rank one subgroups carry the induced Lie subspace topology. Thus one deduces that the embedding of $\operatorname{MBS}(\mathbb{R})_{i}$ into $G(\mathbb{R})$ with respect to the topology $\tau_{\text{KP}}^{\operatorname{as}}$ is continuous. To be more precise, the topology $\tau_{\text{KP}}^{\operatorname{as}}$ comes from the embedding of relative rank one subgroups equipped with the Lie topology and by the closed-subgroup theorem (cf. [HN12, Theorem 9.3.7]) the maximal bounded subgroups carries the same Lie group topology. This means that the topology $\tau_{\text{KP}}^{\operatorname{as}}$ is coarser than $\tau_{\text{MB}}^{\operatorname{as}}$.